OpenStudy (anonymous):

Solve 2 log x = log 36 Answer x = 72 x = 6 x = ±6 x = 18

OpenStudy (across):

\[2\log(x)=\log(36)\]\[\log(x^2)=\log(36)\]\[x^2=36\]\[x=±6\]

jimthompson5910 (jim_thompson5910):

\[\large 2\log(x)=\log(36)\] \[\large \log(x^2)=\log(36)\] \[\large x^2=36\] \[\large x=\sqrt{36}\] \[\large x = 6\] Note: We're only dealing with the positive square root since x must be positive.

jimthompson5910 (jim_thompson5910):

x must be positive because you cannot take the log of a negative number